OpenStudy (anonymous):

x 2/4 can be written as 3 sqrt x^2 2 sqrt x ^ 3 x sqrt x The expression cannot be re-written.

OpenStudy (anonymous):

Can you clarify what "x 2/4" means? Do you mean \[\large x^{2/4}?\]

OpenStudy (anonymous):

yes thats what i mean

OpenStudy (anonymous):

x^{2/4} is the same as x^{1/2}, which is the same as √x

OpenStudy (anonymous):

the answer i got for this was the first one is that correct?

OpenStudy (anonymous):

The first two answer choices have 3's in them and I don't know why.

OpenStudy (anonymous):

thats how they're written on my test

OpenStudy (anonymous):

None of those answer choices look correct. \[\large x^{2/4} = \sqrt[4]{x^2} = \sqrt{x}\]

OpenStudy (anonymous):

thats what i have on there

OpenStudy (anonymous):

opps sorry its supposed to be 2/3